# Filter Like You Test: Data-Driven Data Filtering for CLIP Pretraining

### v1

@@ Abstract
We introduce Filter Like You Test (FLYT), a method for curating large-scale vision-language datasets that learns the usefulness of each data point as a pretraining example. FLYT trains a scoring model that learns to weigh each example using gradient signals from downstream tasks training sets. Using the same training methodology, we develop Mixing-FLYT (M-FLYT), which takes the per-example scores generated by different scoring methods and learns to unify them into a single score. Our training methodology naturally produces a distribution over the training examples, which we leverage through Soft Cap Sampling (SCS), a strategy for obtaining a filtered pretraining dataset from per-example probabilities that samples examples while preventing over-representation through a repetition penalty. Using all three methods, we achieve 40.1% ImageNet zero-shot accuracy on the DataComp medium scale filtering benchmark, a 1.9% absolute accuracy increase over all previous results and a 5.5% increase over results that—like us—use only public resources.
@@ 1 Introduction
Figure 1: In each FLYT training loop, the scoring model takes a batch of upstream data and generates a score for each example. These scores are converted to weights via softmax. A reference model processes the upstream batch, and the resulting embeddings, together with the weights, are used to compute a weighted CLIP loss. The reference model is then updated using gradients from this loss. Next, this updated reference model processes downstream data to compute a downstream loss, which produces a gradient signal that passes through the updated reference model all the way to the scoring model parameters.
@@
We introduce three methods that progressively build upon each other to achieve state-of-the-art performance on the DataComp medium scale filtering benchmark [15]. At the core is Filter Like You Test [^1] The name is inspired by FLYP [16]. (FLYT), a method for training a scoring model that predicts the weight an image-text pair should have in CLIP pretraining. Mixing-FLYT (M-FLYT) applies the same training process to merge existing scoring methods. The third method, Soft Cap Sampling (SCS), is a simple sampling strategy that leverages the probabilistic structure of FLYT’s scores to construct the final pretraining dataset.
@@
FLYT recasts data selection as data re-weighting [34], and learns to upweight (and hence select) the pretraining data that contributes the most to the model’s downstream performance. We jointly train two models: a scoring model that evaluates the usefulness of training examples, and a reference model that provides feedback on these evaluations. During training, the scoring model assigns weights to each example in the batch, which are then used in our weighted CLIP loss function to update the reference model. The updated reference model is then evaluated on downstream data, allowing gradient signal from the downstream loss to flow through it, to the weights, and finally back to the scoring model. This allows the scoring model to give higher weights to examples that reduce the downstream loss of the reference model. We then use the trained scoring model to generate per-example scores. Figure 1 illustrates a single training step of FLYT and Algorithm 1 provides pseudocode. When applying on fixed image/text embeddings, FLYT achieves 31.6% zero-shot ImageNet accuracy, and 32.3% on the average of 38 DataComp evaluation tasks, comparable to the best baseline from Gadre et al. [15].
@@
Previous work aggregates different filtering methods to obtain the best results. They either sum together different scores [20], take the intersection/union of the filtered datasets [15, 27, 20, 46] or pipeline the filtering methods one after the other [49, 46]. Departing from heuristic approaches, we leverage FLYT to directly optimize over a space of mixing methods defined by a simple linear combination of scores. Replacing the aggregation method of [20] with M-FLYT improves ImageNet accuracy by 1.3%. More importantly, M-FLYT can aggregate more scores. When used with 12 different scores, it outperforms the best single score by 1.7% on ImageNet accuracy, and by at least 0.8% on average accuracy.
@@
FLYT produces a probability distribution over the training examples, allowing us to obtain a filtered subset by simply sampling from the distribution. However, sampling with replacement produces too many repetition of the top-scoring examples, and sampling without replacement produces too few. To interpolate between the two sampling methods we propose SCS, a strategy that samples with replacement but reduces repetition using a penalty term (see Algorithm 2). Using only publicly available resources, M-FLYT with SCS has 40.1% ImageNet zero-shot accuracy, which is a 1.9% absolute increase compared to the top entry in the DataComp leaderboard, and a 5.5% improvement compared to the best entry that uses only publicly available resources. Our filtered dataset also achieves 37.5% average accuracy across all evaluation tasks, which is within 1.3% of the best in the leaderboard, and a 0.2% improvement over the best entry using only publicly available resources. [^2] This comparison is made against results reported in the DataComp leaderboard, some of used the full 128M DataComp medium candidate pool, while we were only able to download 119M examples. Table 1 compares our filtered dataset to filtered datasets of prior work when constrained to our downloaded data. On the DataComp small scale (which we did not evaluate until the very end of the project) our method sets a new state-of-the-art in both ImageNet and average accuracy (see Table 3).
Input: Scoring model ${q}_{{\phi}_{0}}$ with initial parameters ${\phi}_{0}$, reference model ${f}_{{\theta}_{0}}$ with initial parameters ${\theta}_{0}$, upstream dataset ${D}$ and downstream dataset ${C}$
Parameters: Number of steps $T$, upstream data batch size ${B}$, downstream data batch size ${B^{\prime}}$, weighted loss ${\ell_{\mathrm{WCL}}}$, and downstream loss ${\ell_{\mathrm{DS}}}$
@@ 7       ${\hat{L}_{\mathrm{down}}}={\ell_{\mathrm{DS}}}({f}_{{\theta}_{t+1}}({{\hat{z}}_{1:{B^{\prime}}}}))$
Input: Training examples ${z}_{1:{M}}$ and their scores ${s}_{1:{M}}$
Parameters: Repetition penalty ${\alpha}$, target dataset size ${N}$ and sample count per iteration ${G}$.
1 ${\hat{D}}=\left[\text{ }\right]$ while $\lvert{\hat{D}}\rvert<{N}$ do
@@ 5       ${s}_{i_{1:{G}}}={s}_{i_{1:{G}}}-{\alpha}$
We share all the code required to reproduce our work, including FLYT, M-FLYT training, and SCS sampling at https://github.com/formll/FLYT. Additionally, we publish our best-performing scoring and mixing models along with the resulting filtered datasets, and the M-FLYT input scores at https://huggingface.co/collections/formll/flyt-67bb167366ec0fa0d5b8e4bd.
@@ 2 Related work
Recognizing the importance of data quality for CLIP training, Gadre et al. [15] introduces the DataComp benchmark. DataComp fixes the data pool as well as the training and evaluation procedures, and allows researchers to focus on comparing filtering methods while keeping all other parameters fixed. Like us, many papers use the DataComp benchmark to evaluate their filtering methods [49, 27, 20, 14, 46, 25]. Below, we survey the literature on data filtering for DataComp.
@@
Gadre et al. [15] consider several baseline filtering methods, and obtain their best results with a combination of CLIP score and a clustering-based filtering method using similarity to the ImageNet training set. T-MARS [27] observe that images containing texts that overlap with the caption received high CLIP scores, while not contributing much to the model training. To address this issue, they mask out these texts before calculating CLIP score. For better results, they take the intersection of this approach and their variation of Maini et al. [26]. Yu et al. [49] develop a sequential pipeline that combines many filtering techniques, which they divide to three approaches: single modality filtering, cross modality filtering, and distribution alignment. DFN [14] use CLIP score from a model fine-tuned on a large high-quality private data. HYPE [20] suggest that relying only on the image-text alignment information produced by CLIP score is insufficient, and propose to also consider the individual information of the images and text. They use hyperbolic embeddings to generate a hyperbolic similarity score in addition to image and text specificity scores, which they sum with the traditional Euclidean CLIP score and the image-based score. NegCLIPLoss [46] introduce a variation to the standard CLIP score that normalizes it by the similarity score of the example with other examples in a batch, and after filtering using this score, they filter the remaining dataset using a downstream examples similarity metric. While these heuristic approaches are successful, our approach aims to directly learn how to estimate an example’s quality.
@@
Given a scoring of the data pool, most filtering approaches select the top 15-20% of examples, resulting in each example appearing 5–7 during training in the DataComp benchmark. Goyal et al. [17] analyzed the quality-quantity trade-off of dataset filtering, and found that for the compute budget of the medium scale DataComp benchmark selecting the top 20% of examples is optimal, with more aggressive filtering leading to diminishing returns. However, their analysis only considers uniform sampling while top performing methods created non-uniformly sampled datasets through concatenation of multiple subsets of data. In their work, HYPE [20] concatenate their selected subset with DFN’s subset, while NegCLIPLoss [46] concatenate their subset with both HYPE and DFN. This implicitly increases the sampling frequency of examples selected by multiple methods, with examples chosen by all three methods being sampled approximately 10 times during training. Our proposed SCS method extends this approach of non-uniform sampling using a simple probabilistic procedure; in our top performing dataset, individual examples appear up to 56 times.
@@ 3 Method
In this section, we describe FLYT in detail. We introduce our notation by going over standard CLIP training (Section 3.1), then move to describing FLYT (Section 3.2), M-FLYT (Section 3.3), and SCS (Section 3.4). Finally, we cover the weighted CLIP loss we use to train FLYT and M-FLYT (Section 3.5).
@@ 3.2 Filter Like You Test (FLYT)
In this section we describe FLYT. Our goal is to learn a scoring model ${q}_{\phi}:{D}\to\mathbb{R}$, parameterized by ${\phi}$, that maps each example ${z}\in{D}$ to a score ${s}$, such that the higher ${s}$ is, the more useful ${z}$ is as a pretraining example. We overload notation to let
${{s}_{1:{B}}}={{q}_{\phi}}({{z}_{1:{B}}})$
@@
${\mathrm{softmax}}({s}_{i})=\frac{\exp({{s}_{i}})}{\sum_{j=1}^{{B}}\exp({{s}_{j}})},$
${{w}_{1:{B}}}={\mathrm{softmax}}({{s}_{1:{B}}}).$
@@
Finally, we generate per-example scores using our trained scoring model and use these scores for data filtering.
3.3 Mixing-FLYT (M-FLYT)
Inspired by past success in mixing multiple filtering heuristics [15, 20, 46, 49], we consider another application of FLYT.
Instead of training a scoring model ${q}:{D}\to\mathbb{R}$ that predicts the per-example scores directly from ${z}$, we take advantage of other scoring methods and train what we call a mixing model. Our mixing model ${p}:\mathbb{R}^{{k}}\to\mathbb{R}$ takes as input ${k}$ different scores, generated by ${k}$ different scoring methods, and outputs a single unified score ${s}$. Other than the different input of the mixing model, we train M-FLYT in the exact same way as FLYT.
Using M-FLYT gives us the advantage of not needing to learn example importance from scratch, and instead we use the information gained by other scoring methods. In fact, we notice that a simple linear mixing model is enough for obtaining strong results.
3.4 Soft Cap Sampling (SCS)
We now consider the task of obtaining a filtered pretraining dataset from per-example scores. A principled data filtering approach should use these scores for deciding the number of repetitions of each example in the filtered dataset. It is not clear how to do that when using a similarity based score as in most prior work. In contrast, a sampling strategy emerges directly from the definition of FLYT, since the scores are naturally interpreted as the learned (log) probability of including an example in the batch.
@@ 3.5 Weighted CLIP loss
    Most importantly, a larger weight means the corresponding example is more represented in the loss.
@@
To realize these principle we modify the standard CLIP loss Equation 1, showing the modifications in red:
@@
${\ell^{\mathrm{w}}_{\text{contrast}}}\left(u_{1:B},v_{1:B},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}{{w}_{1:{B}}}}\right)$
$\qquad=\sum_{i=1}^{{B}}-{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}w_{i}}\log\left(\frac{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}w_{i}}e^{{\tau}\langle u_{i},v_{i}\rangle}}{\sum_{j=1}^{{B}}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}w_{j}}e^{{\tau}\langle u_{i},v_{j}\rangle}}\right).$

### v2

@@ Abstract
We introduce Filter Like You Test (FLYT), an algorithm for curating large-scale vision-language datasets that learns the usefulness of each data point as a pretraining example. FLYT trains a scoring model that learns to weigh each example’s features using gradient signals from downstream tasks training sets. Based on FLYT, we implement Mixing-FLYT (M-FLYT), which takes the per-example scores generated by different scoring methods as features, and learns to unify them into a single score. FLYT naturally produces a distribution over the training examples, which we leverage through Soft Cap Sampling (SCS), a strategy for obtaining a filtered pretraining dataset from per-example probabilities that samples examples while preventing over-representation through a repetition penalty. Using these methods, we achieve 40.1% ImageNet zero-shot accuracy on the DataComp medium scale filtering benchmark, a 2% absolute accuracy increase over all previous results and a 5.5% increase over results that—like us—use only public resources. Our approach also yields 37.7% on the average of 38 DataComp evaluation tasks, outperforming previous public-resource approaches by 0.4%.
@@ 1 Introduction
Figure 1: In each FLYT training loop, the scoring model takes features extracted from a batch of upstream data and generates a score for each example. These scores are converted to weights via softmax. A reference model processes the upstream batch, and the resulting embeddings, together with the weights, are used to compute a weighted CLIP loss. The reference model is then updated using gradients from this loss. Next, this updated reference model processes downstream data to compute a downstream loss, which produces a gradient signal that passes through the updated reference model all the way to the scoring model parameters.
@@
We introduce methods that together achieve state-of-the-art performance on the DataComp medium scale filtering benchmark [17]. At the core is Filter Like You Test [^1] The name is inspired by FLYP [18]. (FLYT), a method for training a scoring model that predicts the weight an image-text pair should have in CLIP pretraining. Mixing-FLYT (M-FLYT) is an implementation of FLYT that uses scores generated by existing scoring methods as input features to the scoring model. Complementing FLYT, Soft Cap Sampling (SCS) is a simple sampling strategy that leverages the probabilistic structure of FLYT’s scores to construct the final pretraining dataset.
@@
FLYT recasts data selection as data re-weighting [37], and learns to upweight (and hence select) the pretraining data that contributes the most to the model’s downstream performance. We jointly train two models: a scoring model that evaluates the usefulness of training examples using extracted features, and a reference model that provides feedback on these evaluations. During training, the scoring model assigns weights to each example in the batch, which are then used in our weighted CLIP loss function to update the reference model. The updated reference model is then evaluated on downstream data, allowing gradient signal from the downstream loss to flow through it, to the weights, and finally back to the scoring model. This allows the scoring model to give higher weights to examples that reduce the downstream loss of the reference model. We then use the trained scoring model to generate per-example scores. Figure 1 illustrates a single training step of FLYT and Algorithm 1 provides pseudocode.
@@
Previous work aggregates different filtering methods to obtain the best results. They either sum together different scores [22], take the intersection/union of the filtered datasets [17, 29, 22, 49] or pipeline the filtering methods one after the other [53, 49]. Departing from heuristic approaches, we leverage FLYT to directly optimize over a space of mixing methods defined by a simple linear combination of scores. Replacing the aggregation method of [22] with M-FLYT improves ImageNet accuracy by 1.3%. More importantly, M-FLYT can aggregate more scores. When used with 12 different scores, it outperforms the best single score by 1.7% on ImageNet accuracy, and by at least 0.8% on average accuracy. M-FLYT outperforms all baseline aggregation methods we evaluated, including summation and ImageNet accuracy-weighted sums by 0.5% on ImageNet accuracy and 0.6% on average accuracy.
A different natural choice for FLYT’s feature extractor component is a multimodal neural network, such as CLIP; we refer to this Embeddings-FLYT (E-FLYT). While E-FLYT does not improve the state-of-the-art on its own, incorporating its scores as an input to M-FLYT leads to improved overall performance.
@@
FLYT produces a probability distribution over the training examples, allowing us to obtain a filtered subset by simply sampling from the distribution. However, sampling with replacement produces too much repetition of the top-scoring examples, and sampling without replacement produces too few. To interpolate between the two sampling methods we propose SCS, a strategy that samples with replacement but reduces repetition using a penalty term (see Algorithm 2). Using only publicly available resources, M-FLYT with SCS has 40.1% ImageNet zero-shot accuracy, which is a 2% absolute increase compared to the best prior work, and a 5.5% improvement compared to the best prior work that uses only publicly available resources. Our filtered dataset also achieves 37.7% average accuracy across all evaluation tasks, which is within 0.1% of our reproduction of the best previously published filtering method, and a 0.4% improvement over the best prior work using only publicly available resources. On the DataComp small scale benchmark (which we did not evaluate until the end of the project) our method outperforms previous work on both ImageNet and average accuracy (see Table 3).
1 Input: Scoring model ${q}_{{\phi}_{0}}$ with initial parameters ${\phi}_{0}$, reference model ${f}_{{\theta}_{0}}$ with initial parameters ${\theta}_{0}$, feature extractor ${\Psi}$, upstream dataset ${D}$, downstream dataset ${C}$
2 Parameters: Number of steps $T$, upstream data batch size ${B}$, downstream data batch size ${B^{\prime}}$, weighted loss ${\ell_{\mathrm{WCL}}}$, and downstream loss ${\ell_{\mathrm{DS}}}$
@@ 10       ${\hat{L}_{\mathrm{down}}}={\ell_{\mathrm{DS}}}({f}_{{\theta}_{t+1}}({{\hat{z}}_{1:{B^{\prime}}}}))$
1 Input: Training examples ${z}_{1:{M}}$ and their scores ${s}_{1:{M}}$
2 Parameters: Repetition penalty ${\alpha}$, target dataset size ${N}$ and sample count per iteration ${G}$
3 ${\hat{D}}=\left[\text{ }\right]$
4 while $\lvert{\hat{D}}\rvert<{N}$ do
@@ 8       ${s}_{i_{1:{G}}}={s}_{i_{1:{G}}}-{\alpha}$
We share all the code required to reproduce our work, including E-FLYT, M-FLYT training, and SCS sampling at https://github.com/formll/FLYT. Additionally, we publish our best-performing scoring and mixing models along with the resulting filtered datasets, and the M-FLYT input scores at https://huggingface.co/collections/formll/flyt-67bb167366ec0fa0d5b8e4bd.
@@ 2 Related work
Recognizing the importance of data quality for CLIP training, Gadre et al. [17] introduces the DataComp benchmark. DataComp fixes the data pool as well as the training and evaluation procedures, and allows researchers to focus on comparing filtering methods while keeping all other parameters fixed. Below, we survey the literature on data filtering for DataComp.
@@
Gadre et al. [17] consider several baseline filtering methods, and achieved best results by combining CLIP score with a clustering-based filtering method using similarity to the ImageNet training set. T-MARS [29] observe that images containing texts that overlap with the caption received high CLIP scores, while not contributing much to the model training. To address this issue, they mask out these texts before calculating CLIP score. For better results, they take the intersection of this approach and their variation of Maini et al. [28]. Yu et al. [53] developed a pipeline merging various techniques categorized as: single modality filtering, cross modality filtering, and distribution alignment. DFN [16] use CLIP score from a model fine-tuned on a large high-quality private data. HYPE [22] argues CLIP’s image-text alignment alone is insufficient, and adds individual information of the images and texts. They use hyperbolic embeddings to generate a hyperbolic similarity score, and image and text specificity scores, which they sum with the traditional Euclidean CLIP and the image-based scores. s-CLIPLoss [49] introduce a variation to the standard CLIP score that normalizes it by the example’s similarity score with other examples in a batch, and after filtering using this score, filters remaining data using a downstream example similarity metric. While these heuristic approaches are successful, our approach directly learn how to estimate an example’s quality.
@@
Given a scoring of the data pool, most filtering approaches select the top 15-20% of examples, resulting in each example appearing 5–7 times during training in the DataComp benchmark. Goyal et al. [19] analyzed the quality-quantity trade-off of dataset filtering, and found that for the compute budget of the medium scale DataComp benchmark selecting the top 20% of examples is optimal, with more aggressive filtering leading to diminishing returns. However, their analysis only considers uniform sampling while top performing methods created non-uniformly sampled datasets through concatenation of multiple subsets of data [22, 49]. Our proposed SCS method extends this approach of non-uniform sampling using a simple probabilistic procedure; in our top performing dataset, individual examples appear up to 56 times.
@@
Concurrent work.
Concurrent with our research, two related works also developed data filtering algorithms evaluated on the DataComp benchmark. Xu et al. [52] introduced EcoDatum which transforms unimodal and multimodal scores into discrete labels and uses a weak supervision ensemble model to aggregate them based on agreements and disagreements [36]. Their approach is similar to M-FLYT in that it combines multiple filtering methods, but unlike us, it doesn’t optimize a downstream objective for aggregation. Separately, Engstrom et al. [14] proposed metagradient descent (MGD), a model-free dataset curation method that, similar to our approach, leverages gradient signals from downstream tasks. While MGD learns to score each example separately, we train a scoring model that applies to unseen examples, which allows us to extend our method to larger dataset scales at a low computational cost. Additionally, whereas MGD optimizes for all available downstream tasks in the DataComp benchmark, we use only the ImageNet training set as the downstream data. MGD sets a new state-of-the-art average accuracy of 40.2% on the DataComp medium benchmark, though its ImageNet accuracy is only 27%.
@@ 3 Method
In this section, we describe FLYT in detail. We introduce our notation by going over standard CLIP training (Section 3.1), then move to describing FLYT (Section 3.2). We then go over different feature extractors for FLYT’s implementation (Section 3.3), followed by our choice of downstream data (Section 3.4), and explaining the weighted CLIP loss we use to train FLYT (Section 3.5). Finally, we present SCS (Section 3.6).
@@ 3.2 Filter Like You Test (FLYT)
In this section we describe FLYT. Our goal is to learn a scoring model ${q}_{\phi}$ parameterized by ${\phi}$. The scoring model takes a feature vector ${\zeta}={\Psi}({z})$ produced by applying feature extractor ${\Psi}$ to example ${z}$, and maps it to a score ${s}$, such that the higher ${s}$ is, the more useful ${z}$ is as a pretraining example. We overload notation to let
${{s}_{1:{B}}}={{q}_{\phi}}({{\zeta}_{1:{B}}})$
@@
${{w}_{1:{B}}}={\mathrm{softmax}}({{s}_{1:{B}}})$
where
${\mathrm{softmax}}({s}_{i})=\frac{\exp({{s}_{i}})}{\sum_{j=1}^{{B}}\exp({{s}_{j}})}.$
@@
Finally, we score each example with our trained scoring model, then filter based on these scores.
3.3 Feature extractors
Mixing-FLYT (M-FLYT)
Inspired by past success in mixing multiple filtering heuristics [17, 22, 49, 53], we implement FLYT using existing scoring methods as our feature extractor. Specifically, the feature extractor ${\Psi}$ produces a vector ${\zeta}\in\mathbb{R}^{{k}}$ of ${k}$ different scores, generated by ${k}$ different scoring methods. Our scoring model ${q}_{\phi}:\mathbb{R}^{{k}}\to\mathbb{R}$ then takes as input the different scores to output a single unified score ${s}$. We refer to this implementation of a scoring model as a mixing model.
Using M-FLYT gives us the advantage of not needing to learn example importance from scratch, and instead use the information gained by other scoring methods. In fact, we notice that a simple linear mixing model is enough for obtaining strong results.
Embedding-FLYT (E-FLYT)
With E-FLYT, we use a pretrained CLIP model as the feature extractor ${\Psi}{}$ to generate image and text embeddings. Our scoring model ${q}_{\phi}$ then maps the concatenated image and text embeddings to a score ${s}$.
Despite CLIP’s ability to produce semantically rich representations, our current E-FLYT implementation does not yet match M-FLYT’s performance. Section C.7 documents our experiments, which show improvements but suggest further work is needed to fully leverage these embeddings.
3.4 Choice of downstream data
FLYT requires high-quality downstream data to effectively learn which examples are more valuable for pretraining. For our implementation, we use the ImageNet training set as a proxy for downstream task data, which aligns with the DataComp competition rules permitting “use of other images associated with these tasks (e.g., supervised training sets)” [17]. This approach follows standard practice, as all other top-performing methods [17, 49, 53, 16] similarly utilize the ImageNet training set for their filtering algorithms. The DataComp benchmark includes 38 diverse tasks, some of which differ significantly from ImageNet. This range of evaluation tasks means that the average downstream score provides a practical measure of model performance on unfamiliar tasks.
@@ 3.5 Weighted CLIP loss
    Crucially, a larger weight means the corresponding example is more represented in the loss.
@@
To realize these principle we modify the CLIP loss Equation 1, showing modifications in red:
@@
${\ell^{\mathrm{w}}_{\text{contrast}}}\left(u_{1:B},v_{1:B},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}{{w}_{1:{B}}}}\right)=\sum_{i=1}^{{B}}-{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}w_{i}}\log\left(\frac{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}w_{i}}e^{{\tau}\langle u_{i},v_{i}\rangle}}{\sum_{j=1}^{{B}}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}w_{j}}e^{{\tau}\langle u_{i},v_{j}\rangle}}\right).$
3.6 Soft Cap Sampling (SCS)
We now consider the task of obtaining a filtered pretraining dataset from per-example scores. A principled data filtering approach should use these scores for deciding the number of repetitions of each example in the filtered dataset. It is not clear how to do that when using a similarity based score as in most prior work. In contrast, a sampling strategy emerges directly from the definition of FLYT, since the scores naturally correspond to the (log) probabilities of including examples in the batch.